OpenStudy (anonymous):

A raffle has a grand prize of a Caribbean cruise valued at $6000 with a second prize of a resort vacation valued at $1400. If each ticket costs $5 and 10400 tickets are sold, what are the expected winnings for a ticket buyer? Express to at least three decimal place accuracy in dollar form (as opposed to cents).

OpenStudy (dumbcow):

expected winnings = sum of (net gain)*(probability) 3 cases: win grand prize, win 2nd, or win nothing grand prize: net gain = 6000 -5 = 5995 probability = 1/10,400 2nd prize: net gain = 1400 -5 = 1395 probability = 1/10,400 win nothing: net gain = -5 probability = 10,398/10,400 --> 5995/10400 + 1395/10400 - (5*10398)/10400 --> -44600/10400 --> -4.29 *odds are definately not in your favor :)

OpenStudy (anonymous):

Thank you so much!!

OpenStudy (dumbcow):

your welcome